OpenStudy (anonymous):

Simplify the radical expression .. r7 + r3

jhonyy9 (jhonyy9):

so what can you factorizing out ?

OpenStudy (anonymous):

I'm not familiar with factoring radicals so I have no idea

jhonyy9 (jhonyy9):

radical expression ??? where is the radical indice ?

OpenStudy (anonymous):

lol \[\sqrt{75} + \sqrt{3}\]

OpenStudy (anonymous):

I think that's what it is and what's a radical indice? o.o

jhonyy9 (jhonyy9):

yes now from these result that there are squareroot of 75 0n2 of 3 so can you rewrite the radical expression in exponent form ?

OpenStudy (anonymous):

uhmm..... -confused-

jhonyy9 (jhonyy9):

so the squareroot has ane xponent form so sqrt(7) = 7^(1/2)

jhonyy9 (jhonyy9):

sqrt(25) = ?

OpenStudy (anonymous):

Where did 7 and 25 come from? e.O

jhonyy9 (jhonyy9):

just for example

OpenStudy (anonymous):

And I apologize, I just realized my question is stated incorrectly

jhonyy9 (jhonyy9):

this is right can you answer my example ?

OpenStudy (anonymous):

square root of 25 is 5 if that's what you mean

jhonyy9 (jhonyy9):

yes how can you rewriting 75 using 25 ?

jhonyy9 (jhonyy9):

like product of

jhonyy9 (jhonyy9):

so 75=25*3

jhonyy9 (jhonyy9):

this mean that sqrt75 = sqrt(25*3) = sqrt25 * sqrt3 = ?

jhonyy9 (jhonyy9):

= 5sqrt3

OpenStudy (anonymous):

\[5\sqrt{3} ?\]

OpenStudy (anonymous):

oohh I see

jhonyy9 (jhonyy9):

and the result ?

OpenStudy (anonymous):

There's another step?? o:

jhonyy9 (jhonyy9):

no just assum these

OpenStudy (anonymous):

>.<

jhonyy9 (jhonyy9):

sqrt75 + sqrt3 = 5sqrt3 +sqrt3 = ?

OpenStudy (anonymous):

\[\sqrt{3 + 5} \sqrt{5}\]

jhonyy9 (jhonyy9):

no no

OpenStudy (anonymous):

dang it

jhonyy9 (jhonyy9):

sqrt(x) +sqrt(x) = ?

OpenStudy (anonymous):

x? I have no idea

jhonyy9 (jhonyy9):

this is just an example

OpenStudy (anonymous):

I know.. and I'm failing at it

jhonyy9 (jhonyy9):

so a+a= ?

OpenStudy (anonymous):

\[a ^{2}\] or a?

jhonyy9 (jhonyy9):

no you have confused these all

jhonyy9 (jhonyy9):

1+1= ?

OpenStudy (anonymous):

lmaooo omg

OpenStudy (anonymous):

2

jhonyy9 (jhonyy9):

so than a+a= ?

OpenStudy (anonymous):

2a....?

jhonyy9 (jhonyy9):

yes hurry up

jhonyy9 (jhonyy9):

and now sqrta +sqrta = ?

OpenStudy (anonymous):

2sqrta?

jhonyy9 (jhonyy9):

yeahhh

jhonyy9 (jhonyy9):

so than 5sqrt3 + sqr3 = ?

OpenStudy (anonymous):

\[6\sqrt{6}\] ... or \[6 \sqrt{3+3}\]

jhonyy9 (jhonyy9):

why ???? 2sqrta + 3sqrta = ?

OpenStudy (anonymous):

xD I suck....! pffft uhmmm 5sqrta

jhonyy9 (jhonyy9):

yes so than 5sqrt3 + sqrt3 = ?

OpenStudy (anonymous):

See... that confuses me... there's nothing in front of the sqrt3 so is it 1 or 0? and am I supposed to add the sqrt3's together or what?

jhonyy9 (jhonyy9):

yes 5sqrt3 + 1sqrt3 =

OpenStudy (anonymous):

6sqrt3 ?

jhonyy9 (jhonyy9):

yes so thia will be the right answer

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

OMGEE THANK YOU ! ughhhhhh that took a lifetime, thank you so much for enduring this with me >.<

jhonyy9 (jhonyy9):

was my pleasure good luck bye